OpenStudy (anonymous):

PLEASE HELP 2x^2+4=-5-2x^2 or ( the 5 is a positive) 2x^2+4=5-2x^2 solve using the quadratic formula Now I did this probably 5 times over and havnt gotten the right answer will show my work below.

OpenStudy (anonymous):

first get everything on the same side

OpenStudy (anonymous):

\[-4\pm \sqrt{4^{2}-4(4)(-5)} all \div by 2(4)\]

OpenStudy (anonymous):

4x^2 + 0x + 9 = 0 is your new equation now a=4 , b=0, c= 9

OpenStudy (anonymous):

\[or -4\pm \sqrt{4^{2}-4(4)(5)} all \div by 2(4)\]

OpenStudy (anonymous):

what type the quadratic formula first using letters and you will know what your mistake is

OpenStudy (anonymous):

OHHHH there is no x i am so sorry i overlooked that simple part

OpenStudy (anonymous):

sorry I meant to say type** not what type

OpenStudy (anonymous):

haha no reason to be sorry so do you want to show what the new formula looks like?

OpenStudy (anonymous):

yes \[-0\pm \sqrt{0^{2}-4(4)(9)} all diveded by 2(4) \]

OpenStudy (anonymous):

let me do it quickly and see if i get an answer because now I think i can do it

OpenStudy (anonymous):

yes that should work you know imaginary numbers right?

OpenStudy (anonymous):

no what are those and I cant do this because the square root has a -144 what is wrong with that ?

OpenStudy (anonymous):

oh wait why does it say or the 5 is a positive at the beginning of your question?

OpenStudy (anonymous):

because the question where online and I looked really close and it looked like a -5, but when i did it my way that didn't work so I tried positive and got a closer answer but if -5 works then it is that

OpenStudy (anonymous):

If the five is a positive the c = -1 and you will have a real answer and if the 5 is a - then it is -9 and your answer will be i*sqrt(144) the i stands for sqrt of -1

OpenStudy (anonymous):

well it was multiple choice the choices are a) \[1+\sqrt{6} \div4 and 1-\sqrt{6}\]

OpenStudy (anonymous):

b) \[-1+\sqrt{6} and -1-\sqrt{6}\]

OpenStudy (anonymous):

c) \[1+\sqrt{6} \div4 and 1-\sqrt{6}\div4\]

OpenStudy (anonymous):

d) \[1+\sqrt{6} and 1-\sqrt{6}\]

OpenStudy (anonymous):

and if i say divide its the line under the numbers with the 4 on the bottom

OpenStudy (anonymous):

sorry I shouldn't have said you answer I mean the part under the sqrt root you answer would be -i12/8 = + or - (-i3/2) or if 5 is positive okay wait if none of the answers have an i then the 5 be positive

OpenStudy (anonymous):

oh my you are gona kill me the numbers from the start are wrong i typed them wrong they are \[2x^{2}+4x=-5-2x ^{2} \]

OpenStudy (anonymous):

okay if you equation is 2x^2+4=5-2x^2 then you get 4x^2 - 1 = 0 and that mean a = 4, b = 0 and c = -1 \[(-0 \pm \sqrt{0 - 4*4*-1})/(2*4)\] so you should have \[(\pm \sqrt{16})/(8)\]

OpenStudy (anonymous):

lol it's cool just one second we will figure this out

OpenStudy (anonymous):

its really driving me insane if you could see my notes i tried it every single way and it doesn't give me any of the answers

OpenStudy (anonymous):

so if the 5 is actually a positive the end answer would be \[-4 \pm \sqrt{16x6} / 8 \] \[-4\pm 4\sqrt{6} / 8\]

OpenStudy (anonymous):

and if the 5 is a - then \[-4 \pm \sqrt{64} / 8 \] and we didnt learn that there can be a negative in the square root so this is simply not an answer

OpenStudy (anonymous):

and if we wanted to make things more fun the +5 could also be \[-2 \pm 2 \sqrt{6} /4 \] and this is so very close to the answers but now if you go on you get \[0\sqrt{6} / 4 and -\sqrt{6}\]

OpenStudy (anonymous):

2x^2+4x=5−2x^2 okay lets just assume the 5 starts off positive 4x^2 + 4x + 5 = 0 a = 4, b = 4, c=-5 \[ (−4\pm√(16 - 4*4*-5) )/(2∗4)\] \[ (−4\pm√96 )/(8)\] and I just checked on wolfram with the equation you gave me and it appears to be right and when it is simplified it looks more like one of the choices. (1/2)(-1-sqrt(6))

OpenStudy (anonymous):

how did u get one of the answers since \[-4 + 4 \sqrt{6} /8 \] \[-1 + 1 \sqrt{6} \] \[0 \sqrt{6} / 4\]

OpenStudy (anonymous):

or do you multiply the numbers from the front of the square root ?

OpenStudy (anonymous):

Do you know how to simplify to get to that answer? (−4±√96)/(8) is the same as (−4±√(6*16))/(8) and we can rewrite that as (−4±4√6)/(8) then factor out the four on top to get 4(−1±√6)/(8) and then then 4/8 becomes 1/2 so you have (−1±√6)/2

OpenStudy (anonymous):

thank you so much I understand now so sorry for talking your time But thank you oh so very much

OpenStudy (anonymous):

It's cool remember to read the question very carefully I always read stuff to fast and it messes me up too.